Question 2: Basic Probability (New) After graduation, you accept a position at an automotive manufacturer. Your company just received a shipment of raw materials made of aluminum, stainless steel, and cast iron in three separate boxes A, B and C. Box A contains the aluminum parts, box B contains the stainless steel parts, and box C contains the cast iron parts. Box A contains a total of 100 parts, among which 20 are defective. Box B contains 150 parts, among which 25 are defective. Box C contains 200 parts, among which 15 are defective. Use this information to answer question 2a.
One of the operators in the plant mistakenly combined all three boxes. That means now you have a single box containing all the materials (450 materials in total). Answer question 2b-2d based on this scenario:
Question 2d: You randomly select a part, determine what type of part it is and whether it is defective, place it back in the box and repeat this process so that you have selected a total of two parts from the box. Assuming the trials are independent, what is the probability of getting non-defective stainless steel part both times (round to 3 decimals)?
Answer #1

P(non defective)=P(from A and non defective)+P(from B and non defective)+P(from C and non defective)

=(100/450)*(80/100)+(150/450)*(125/150)+(200/450)*(185/200)=390/450

hence 2d) P(both are non defective) =(390/450)*(390/450)=0.751